# Enhancing AI System Resiliency: Formulation and Guarantee for LSTM Resilience Based on Control Theory

### v4

@@ Abstract
This paper proposes a novel theoretical framework for guaranteeing and evaluating the resilience of long short-term memory (LSTM) networks in control systems. We introduce recovery time as a new metric of resilience in order to quantify the time required for an LSTM to return to its normal state after anomalous inputs. By mathematically refining incremental input-to-state stability ($\delta$ISS) theory for LSTM, we derive a practical data-independent upper bound on recovery time. This upper bound gives us resilience-aware training. Experimental validation on simple models demonstrates the effectiveness of our resilience estimation and control methods, enhancing a foundation for rigorous quality assurance in safety-critical AI applications.
@@ 1.2 Objective
The key tool to achieve our goals is the incremental input-to-state stability ($\delta$ISS) (Angeli,, 2002; Bayer et al.,, 2013), because $\delta$ISS guarantees the resilience of LSTM. Specifically, even if temporary perturbations in the input sequence disrupt inference, the state variables can converge to the same inference values that would have been obtained without perturbations. Sufficient conditions for the LSTM to exhibit the $\delta$ISS properties have already been derived as inequalities of parameters (Terzi et al.,, 2021; Bonassi et al.,, 2023). However, it is not clarified in these studies how long it takes for state variables to converge. We will refine and improve upon these previous studies to achieve the above three goals.
@@ 1.3 Contribution
  • Improvement of sufficient conditions for LSTM to be $\delta$ISS (Section 4).
@@ 2.1 Control System and Stability Theory
In control theory, stability and resilience are important properties. Furthermore, it is well-established that the stability of controlled objects fundamentally determines the overall stability and resilience of control systems. For instance, systems possessing input-to-state stability (ISS) properties achieve global asymptotic stability under closed-loop control (Jiang and Wang,, 2001). Moreover, systems with $\delta$ISS characteristics acquire robustness when implemented with model predictive control (MPC) (Bayer et al.,, 2013).
@@ 2.2 Theoretical LSTM Stability Assurance
Theoretical approaches have investigated sufficient conditions for LSTM networks to satisfy various stability properties. For single-layer LSTMs, conditions for global asymptotic stability (Deka et al.,, 2019), ISS (Bonassi et al.,, 2020), and $\delta$ISS (Terzi et al.,, 2021) have been successively established. Research on sufficient conditions for $\delta$ISS has been extended to multilayer LSTM architectures (Bonassi et al.,, 2023).
@@ 3 Preliminary
In this section, we introduce essential mathematical preliminaries, an important stability concept: $\delta$ISS, and the LSTM.
@@ 3.1 Notation
For a vector $v\in\mathbb{R}^{n}$, $\|v\|=\|v\|_{2}$ represents the Euclidean norm of $v$, and $\|v\|_{\infty}=\max_{i\in[n]}|v_{(i)}|$ represents the maximum norm, where $v_{(i)}$ is the $i$-th component of $v$ and $[n]=\{1,2,\ldots,n\}$. For a discrete-time sequence of vector $v(k),k\in\mathbb{Z}_{\geq 0}$, we denote $v(k_{1}:k_{2})=[v(k_{1}),\ldots,v(k_{2})]$, and $\|v(k_{1}:k_{2})\|_{2,\infty}=\|(\|v(k_{1})\|_{2},\ldots,\|v(k_{2})\|_{2})^{\top}\|_{\infty}$.
For a matrix $A\in\mathbb{R}^{m\times n}$, $\|A\|=\|A\|_{2}$ represents the induced 2-norm of $A$, and $\|A\|_{\infty}$ does the induced $\infty$-norm, that is, $\|A\|_{\infty}=\max_{i}\sum_{j}A_{(i,j)}$. In addition, $\left\lvert{A}\right\rvert$ denotes a matrix whose components are the absolute values of $A$, that is, $\left\lvert{A}\right\rvert\coloneqq(|A_{(i,j)}|)$. For a matrix $A\in\mathbb{R}^{n\times n}$, $\rho(A)$ denotes the spectral radius of $A$, that is, $\rho(A)\coloneqq\max_{i\in[n]}|\lambda_{i}|$, where $\lambda_{i}$ are the eigenvalues of $A$.
@@ 3.2 Stability
Definition 3.2 ($\delta$ISS, Angeli, 2002; Bayer et al., 2013).
System (1) is called incrementally input-to-state stable ($\delta$ISS) if there exist continuous functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that, for any initial states $s_{1}(0),s_{2}(0)\in\mathbb{R}^{n_{s}}$ and for any input sequence $x_{1}(0:t)$, $x_{2}(0:t)\in[-x_{\mathrm{max}},x_{\mathrm{max}}]^{n_{x}\times(t+1)}$, $s_{1}(t)$ and $s_{2}(t)$ satisfy
@@ 3.3 LSTM Network
$x^{(l)}(t)$ =$=\begin{cases}x&\text{for }l=1,\\ h^{(l-1)}(t+1)&\text{for }l=2,\ldots L,\end{cases}$ (3e)
@@
4 $\delta$ISS of LSTM
In this section, we relax the existing $\delta$ISS condition of the LSTM (Terzi et al.,, 2021) by refining the invariant set of the LSTM dynamics. This relaxation contributes to an accurate evaluation of the recovery time introduced in Section 5.
@@ 4.1 Invariant Set in LSTM
4.2 $\delta$ISS of LSTM
Using Proposition 4.1, we provide the following sufficient condition of $\delta$ISS:
Theorem 4.2 ($\delta$ISS of LSTM).
Let $L\in\mathbb{N}$ and $k\in\mathbb{Z}_{\geq 0}$. We assume that $s^{(l)}(0)\in\mathcal{S}^{(l)}(k)$ for all $l\in[L]$. The LSTM is $\delta$ISS if $\rho(A^{{(l)}}_{s}(k))<1$ for all $l\in[L]$, where
@@ 5.1 Definition
Evaluating the recovery time $T_{R}$ is difficult because it depends on the input data, then we utilize the function $\beta$ of $\delta$ISS in Definition 3.2.
@@ 5.2 Upper Bound of Recovery Time
×$\qquad\times\max_{i=l,\cdots,L}\rho(A_{s}^{(i)}(k))^{t}\left(\prod^{L}_{i=l+1}\|a_{x}^{(i)}(k)\|\right)s^{(l)}.$ (23)
@@ 6.1 Estimate
$=\min\{t_{1}\geq 0\ :\ \tilde{\beta}(2\bar{s}(k),t;k)\leq e/\|U_{y}\|\ \text{for any}\ t\geq t_{1}\}$
@@ 7.1.2 Results and Analysis
We calculated $T_{R}$ and $\bar{T}_{R}(k)$ $(k=0,1,\cdots,20)$ for randomly initialized 202020 models, deriving estimated average test errors and correlation coefficients. As shown in Figure 1, the estimation test error consistently decreases as $k$ increases, eventually converging at around $k=20$ steps. Furthermore, the correlation coefficient exhibits a generally increasing (except for a temporary decrease) trend as $k$ increases, which indicates a strong relationship between the true and our estimated values.
Figure 2 compares the estimation accuracy for $k=0$ and $k=20$, demonstrating improved accuracy with a larger $k$. The left side of the red line is the area where $\delta$ISS is guaranteed by Theorem 4.2. Increasing $k$ from 0 to 202020 decreases $\rho(A(k))$ by 32% on average, then all data points are included in the $\delta$ISS guaranteed area.
@@
These results offer empirical support for our theoretical statements in Theorem 5.3 and Remark 5.4: increasing $k$ not only relaxes the sufficient condition of $\delta$ISS but also enhances the accuracy of recovery time estimation.
Figure 1: k-dependency of $\bar{T}_{R}(k)$. Each point represents a pair of the correlation coefficient and estimation error between $\bar{T}_{R}(k)$ and $T_{R}$ for $k=0,1,\ldots,20$ (averaged over 202020 models). Edges connect points corresponding to consecutive $k$ values. The figure shows that, as $k$ increases, the correlation coefficient tends to increase, and the estimation error monotonically decreases.
Figure 2: Accuracy of recovery time estimation. Gray dots represent $T_{R}$ for each model. Orange and blue dots indicate $\bar{T}_{R}(k)$ in $k=0$ and $k=20$ for each model. The red line represents the $\delta$ISS sufficient condition boundary. The models in the left side area of the red line are guaranteed $\delta$ISS by Theorem 4.2. Increasing $k$ from 0 to 202020 decreases $\rho(A(k))$ by 32% on average over 20 models. This figure indicates that increasing $k$ guarantees more models to be $\delta$ISS.
@@ 7.2.2 Results and Analysis
Next, concerning (B), we observed a trade-off between recovery times and MAE loss when comparing the baseline model and our models, and when varying $\varepsilon$ of our model, as illustrated in Figure 4. Specifically, as $\varepsilon$ increases from 0.10.10.1 to 0.50.50.5, we observe a trend of increasing MAE loss and decreasing recovery time. This trend aligns with our theoretical predictions in Section 6.2.
@@
Figure 4: Trade-off controlled by ${\bm{\varepsilon}}$. The red dot represents the baseline value. For each $\varepsilon$ from 0.10.10.1 to 0.50.50.5 in 0.050.050.05 increments value, gray, orange, and blue dots represent MAE and $T_{R},\bar{T}_{R}(0)$, and $\bar{T}_{R}(20)$ pairs, respectively, where MAE is the loss calculated on test data. The figure shows a trade-off between recovery times and MAE when comparing the baseline and our models, and when varying $\varepsilon$ of our model.
@@ 8 Discussion
  • As tasks become increasingly dependent on long-term memory, the decline in the accuracy of recovery time estimation and the accuracy degradation caused by the constraints for resilience are likely to emerge as significant challenges. Relaxing the $\delta$ISS sufficient conditions could be a key approach to address these challenges. This is because, as demonstrated in Theorem 5.3, the $\delta$ISS conditions form the foundation for the recovery time estimation, and relaxing these conditions would contribute to weakening model constraints and mitigating inference accuracy degradation.
@@ Appendix A Other Related Work
In this paper, we define recovery time as the duration from when different input sequences align until the output difference falls below a threshold. We prove that LSTMs with $\delta$ISS properties can effectively estimate this time.
@@
However, deriving sufficient FTS conditions involves strict model constraints, established only for specific models like Hopfield networks (Michalak and Nowakowski,, 2017) and Clifford-valued RNNs (Shen and Meng,, 2023), but not for conventional RNNs or LSTMs. While achieving FTS for LSTMs would theoretically allow for precise recovery time evaluation, we argue that from a practical standpoint, considering numerical errors and rounding, it suffices for output differences to fall below a threshold rather than strictly reaching equilibrium. Thus, we opted for $\delta$ISS instead of FTS.
@@ Appendix B Proofs
$0=\phi\left(0\times 2\right)\frac{1}{2}\leq\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(k)\bar{\phi}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}\right)\overline{\sigma}^{(l)}_{o}(k)=\eta(k).$ (38)
@@
In the following, we show these sequences are monotonically decreasing by induction, therefore converging.
@@
0 0 $\leq\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(0)\overline{\phi}^{(l)}_{c}(0)}{1-\overline{\sigma}^{(l)}_{f}(0)}\right)<1=\eta^{(l)}(-1),$ (39)
@@
$=\overline{\sigma}^{(l)}_{f}(k)\frac{\overline{\sigma}^{(l)}_{i}(k)\overline{\phi}^{(l)}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}+\overline{\sigma}_{i}(k)\overline{\phi}_{c}(k)$ (60)
@@ B.2 Approximation of $\eta^{(l)}(\infty)$
Denote the limits of sequences $\{\overline{\sigma}^{(l)}_{*}(k)\}_{k=0}^{\infty}$ (* denotes $f,i$, or $o$), $\{\overline{\phi}^{(l)}_{c}(k)\}_{k=0}^{\infty}$, and $\{\eta^{(l)}(k)\}_{k=-1}^{\infty}$ as $\overline{\sigma}^{(l)}_{*}(\infty)$, $\overline{\phi}^{(l)}_{c}(\infty)$ and $\eta^{(l)}(\infty)$, respectively. Since $\overline{\sigma}^{(l)}_{*}(\infty)=\sigma(G^{(l)}_{*}(\eta^{(l)}(\infty)))$ and $\overline{\phi}^{(l)}_{c}(\infty)=\phi(G^{(l)}_{c}(\eta^{(l)}(\infty)))$, we have
@@
If the equation (64) could be solved, that is, $\overline{\sigma}^{(l)}_{*}(\infty)$ and $\overline{\phi}^{(l)}_{c}(\infty)$ could be obtained, we would find the loosest sufficient condition for the LSTM to be $\delta$ISS, and the optimal estimate of recovery time. However, it is difficult to solve it analytically, so we approximate it using properties of convex functions.
@@
+[$\quad+\left[\left\{(1+(r^{(l)})^{2t})(\lambda_{1}^{(l)})^{2}+(\nu^{(l)})^{2}\sum_{\tau=1}^{t}(r^{(l)})^{\tau-1}\sum_{\upsilon=1}^{t}(r^{(l)})^{\upsilon-1}\right\}^{2}-4(r^{(l)})^{2t}(\lambda_{1}^{(l)})^{4}\right]^{\frac{1}{2}}(\lambda_{1}^{(l)})^{2(t-1)}$ (126)
@@
Theorem B.5 ($\delta$ISS of LSTM).
Let $L\in\mathbb{N}$ and $k\in\mathbb{Z}_{\geq 0}$. We assume that $s^{(l)}(0)\in\mathcal{S}^{(l)}(k)$ for all $l\in[L]$. The LSTM is $\delta$ISS if $\rho(A^{{(l)}}_{s}(k))<1$ for all $l\in[L]$, where
@@
+$\quad+\left\{\sigma\left(W^{{(l)}}_{f}x_{1}^{{(l)}}(t)+U^{{(l)}}_{f}h_{1}^{{(l)}}(t)+b^{{(l)}}_{f}\right)-\sigma\left(W^{{(l)}}_{f}x_{2}^{{(l)}}(t)+U^{{(l)}}_{f}h_{2}^{{(l)}}(t)+b^{{(l)}}_{f}\right)\right\}\odot c_{2}^{(l)}(t)$ (134)
@@
+$\quad+\left\{\sigma\left(W^{{(l)}}_{i}x_{1}^{{(l)}}(t)+U^{{(l)}}_{i}h_{1}^{{(l)}}(t)+b^{{(l)}}_{i}\right)-\sigma\left(W^{{(l)}}_{i}x_{2}^{{(l)}}(t)+U^{{(l)}}_{i}h_{2}^{{(l)}}(t)+b^{{(l)}}_{i}\right)\right\}$ (137)
@@
+$\quad+\left\{\sigma\left(W^{{(l)}}_{o}x_{1}^{{(l)}}(t)+U^{{(l)}}_{o}h_{1}^{{(l)}}(t)+b^{{(l)}}_{o}\right)-\sigma\left(W^{{(l)}}_{o}x_{2}^{{(l)}}(t)+U^{{(l)}}_{o}h_{2}^{{(l)}}(t)+b^{{(l)}}_{o}\right)\right\}\odot\phi\left(c_{2}^{(l)}(t+1)\right).$ (141)
Since the Lipschitz constant of $\sigma$ is 1/41/41 / 4, we obtain that
@@
for multi-layer LSTM. Thus, LSTM is $\delta$ISS if $\rho(A_{s}^{(l)}(k))<1$ for all $l\in[L]$.
@@
$\sup_{s(0)\in\prod^{L}_{l=1}S^{(l)}_{k}}T_{R}(x,\widehat{x},s(0);e)\leq\min\{t_{1}\geq 0:\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t;k)\leq e/\|U_{y}\|\ \text{for any}\ t\geq t_{1}\},$ (163)
@@ C.1.2 Evaluation
where $t=0,1,2,\ldots,99$. $T_{R}(x,\widehat{x})$ and $\bar{T}_{R}(k)$ were calculated according to Algorithm 1 and Algorithm 2 where $\mathcal{T}=100$. It should be noted that in cases where the model exhibits instability or $\rho(A_{s}(k))\geq 1$, this algorithm computes $T_{R}=\infty$ or $\bar{T}_{R}(k)=\infty$. Calculating statistical measures such as estimation errors or correlation coefficients using infinite values would result in arbitrary and meaningless numbers. Indeed, for some values of $k$ and specific models, our experiments showed that $\bar{T}_{R}(k)=\infty$ and $\rho(A_{s}(k))\geq 1$ occurred. This result is theoretically correct and does not indicate any deficiency in the theory. However, using infinite values to calculate statistical measures such as estimation errors or correlation coefficients would result in arbitrary and meaningless numbers. To address this issue, we replaced infinity values with finite proxies: $\bar{T}_{R}(k)=\mathcal{T}=100$ when computing these statistical measures.
Algorithm 1 Computation of $T_{R}(x,\widehat{x})$

### v5

@@ Abstract
[…]
@@ 1.2 Objective
[…]
@@ 1.3 Contribution
[…]
@@ 2.1 Control System and Stability Theory
[…]
@@ 2.2 Theoretical LSTM Stability Assurance
[…]
@@ 3 Preliminary
[…]
@@ 3.1 Notation
[…]
@@ 3.2 Stability
[…]
@@ 3.3 LSTM Network
$x^{(l)}(t)$ $=\begin{cases}x&\text{for }l=1,\\ h^{(l-1)}(t+1)&\text{for }l=2,\ldots L,\end{cases}$ (3e)
@@
[…]
@@ 4.1 Invariant Set in LSTM
[…]
@@ 5.1 Definition
[…]
@@ 5.2 Upper Bound of Recovery Time
$\qquad\times\max_{i=l,\cdots,L}\rho(A_{s}^{(i)}(k))^{t}\left(\prod^{L}_{i=l+1}\|a_{x}^{(i)}(k)\|\right)s^{(l)}.$ (23)
@@ 6.1 Estimate
$=\min\{t\geq 0:\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t^{\prime};k)\leq e/\|U_{y}\|\ \text{for any}\ t^{\prime}\geq t\}$
@@ 7.1.2 Results and Analysis
We calculated $T_{R}$ and $\bar{T}_{R}(k)$ $(k=0,1,\cdots,20)$ for randomly initialized $20$ models, deriving estimated average test errors and correlation coefficients. As shown in Figure 1, the estimation test error consistently decreases as $k$ increases, eventually converging at around $k=20$ steps. Furthermore, the correlation coefficient exhibits a generally increasing (except for a temporary decrease) trend as $k$ increases, which indicates a strong relationship between the true and our estimated values.
[…]
@@
[…]
Figure 1: k-dependency of $\bar{T}_{R}(k)$. Each point represents a pair of the correlation coefficient and estimation error between $\bar{T}_{R}(k)$ and $T_{R}$ for $k=0,1,\ldots,20$ (averaged over 2020 models). Edges connect points corresponding to consecutive $k$ values. The figure shows that, as $k$ increases, the correlation coefficient tends to increase, and the estimation error monotonically decreases.
[…]
@@ 7.2.2 Results and Analysis
Next, concerning (B), we observed a trade-off between recovery times and MAE loss when comparing the baseline model and our models, and when varying $\varepsilon$ of our model, as illustrated in Figure 4. Specifically, as $\varepsilon$ increases from $0.1$ to $0.5$, we observe a trend of increasing MAE loss and decreasing recovery time. This trend aligns with our theoretical predictions in Section 6.2.
@@
Figure 4: Trade-off controlled by $\bm{\varepsilon}$}. The red dot represents the baseline value. For each $\varepsilon$ from $0.1$ to $0.5$ in $0.05$ increments value, gray, orange, and blue dots represent MAE and $T_{R},\bar{T}_{R}(0)$, and $\bar{T}_{R}(20)$ pairs, respectively, where MAE is the loss calculated on test data. The figure shows a trade-off between recovery times and MAE when comparing the baseline and our models, and when varying $\varepsilon$ of our model.
@@ 8 Discussion
[…]
@@ Appendix A Other Related Work
[…]
@@
[…]
@@ Appendix B Proofs
$0=\phi\left(0\times 2\right)\frac{1}{2}\leq\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(k)\bar{\phi}^{(l)}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}\right)\overline{\sigma}^{(l)}_{o}(k)=\eta^{(l)}(k).$ (38)
@@
In the following, we show that these sequences are monotonically decreasing by induction, therefore converging.
@@
$0$ $\leq\phi\left(\frac{\overline{\sigma}^{(l)}_{i}(0)\overline{\phi}^{(l)}_{c}(0)}{1-\overline{\sigma}^{(l)}_{f}(0)}\right)<1=\eta^{(l)}(-1),$ (39)
@@
$=\overline{\sigma}^{(l)}_{f}(k)\frac{\overline{\sigma}^{(l)}_{i}(k)\overline{\phi}^{(l)}_{c}(k)}{1-\overline{\sigma}^{(l)}_{f}(k)}+\overline{\sigma}^{(l)}_{i}(k)\overline{\phi}^{(l)}_{c}(k)$ (60)
@@ B.2 Approximation of $\eta^{(l)}(\infty)$
[…]
@@
[…]
@@
$\quad+\left[\left\{(1+(r^{(l)})^{2t})(\lambda_{1}^{(l)})^{2}+(\nu^{(l)})^{2}\sum_{\tau=1}^{t}(r^{(l)})^{\tau-1}\sum_{\upsilon=1}^{t}(r^{(l)})^{\upsilon-1}\right\}^{2}-4(r^{(l)})^{2t}(\lambda_{1}^{(l)})^{4}\right]^{\frac{1}{2}}(\lambda_{1}^{(l)})^{2(t-1)}$ (126)
@@
[…]
@@
$\quad+\left\{\sigma\left(W^{{(l)}}_{f}x_{1}^{{(l)}}(t)+U^{{(l)}}_{f}h_{1}^{{(l)}}(t)+b^{{(l)}}_{f}\right)-\sigma\left(W^{{(l)}}_{f}x_{2}^{{(l)}}(t)+U^{{(l)}}_{f}h_{2}^{{(l)}}(t)+b^{{(l)}}_{f}\right)\right\}\odot c_{2}^{(l)}(t)$ (134)
@@
$\quad+\left\{\sigma\left(W^{{(l)}}_{i}x_{1}^{{(l)}}(t)+U^{{(l)}}_{i}h_{1}^{{(l)}}(t)+b^{{(l)}}_{i}\right)-\sigma\left(W^{{(l)}}_{i}x_{2}^{{(l)}}(t)+U^{{(l)}}_{i}h_{2}^{{(l)}}(t)+b^{{(l)}}_{i}\right)\right\}$ (137)
@@
$\quad+\left\{\sigma\left(W^{{(l)}}_{o}x_{1}^{{(l)}}(t)+U^{{(l)}}_{o}h_{1}^{{(l)}}(t)+b^{{(l)}}_{o}\right)-\sigma\left(W^{{(l)}}_{o}x_{2}^{{(l)}}(t)+U^{{(l)}}_{o}h_{2}^{{(l)}}(t)+b^{{(l)}}_{o}\right)\right\}\odot\phi\left(c_{2}^{(l)}(t+1)\right).$ (141)
Since the Lipschitz constant of $\sigma$ is $1/4$, we obtain that
@@
[…]
@@
$\sup_{s(0)\in\prod^{L}_{l=1}S^{(l)}_{k}}T_{R}(x,\widehat{x},s(0);e)\leq\min\{t\geq 0:\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t^{\prime};k)\leq e/\|U_{y}\|\ \text{for any}\ t^{\prime}\geq t\},$